Represented by 3 consecutive odd integers, such that the product of the first and third was equal to 4 less than 5 times the second.
Answer by jorel555(1290)

Let x be your first odd integer; then the second and third ones are x+2 and x+4, respectively:
x(x+4)=5(x+2)-4
x^2+4x=5x+6
x^2-x-6=0
(x-3)(x+2)=0
x=3,-2
The integers are 3,5, and 7!!!!!!!